# Concurrent Composition for Differentially Private Continual Mechanisms

### v2

@@ A.1. Preliminaries
implying that each $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ forms a valid PMF.
A.2.4. Proof of Lemma A.1.
By definition of $\mathrm{RR}$ and of $\mathcal{P}$, for each $t\in\mathbb{N}$ and transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the probability of $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ returning the answers $a_{1},\dots,a_{t}$ to the queries $q_{1},\dots,q_{t}$ is
@@
Thus, the views of any adversary interacting with $\mathcal{M}(s_{b})$ and $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ are identically distributed, implying that these mechanisms are equivalent.
A.3. Construction of $\phi_{t}$
In this section, we formalize the construction of [18] differently to remove the finiteness assumption on $A$.
Roughly speaking, Lyu [18] initializes the values of $\phi_{t}$ with the ones from $\mathrm{L}_{t}$. Then, for each $q_{1},\dots,a_{t-1},q_{t}$, they iterate over all possible answers $a_{t}\in A$ in a fixed order and gradually increase the value of $\phi_{t}(q_{1},\dots,a_{t})$ to satisfy conditions (II) and (III), while ensuring that condition (I) remains valid. This iterative procedure requires $A$ to be finite; otherwise, the construction algorithm does not terminate. To extend this approach to infinite $A$, we define $\phi_{t}$ mathematically and apply the transfinite recursion theorem to prove its existence.
For each $t\in\mathbb{N}$, knowing that $\phi_{t-1}$ exists and satisfies condition (I), we construct $\phi_{t}$ and show it satisfies conditions (I), (II), and (III). For $t=0$, we set $\phi_{0}()=(1,1)$. By Corollary A.16, $\phi_{0}$ satisfies condition (I).
Remark A.22.
In the rest of this section, we assume $\phi_{t-1}$ exists and satisfies condition (I).
In Section A.3.1, we introduce the necessary preliminaries for formally defining $\phi_{t}$ when $A$ is infinite. In Section A.3.2, we construct $\phi_{t}$, and in Section A.3.3, we prove that $\phi_{t}$ satisfies conditions (I) to (III).
A.3.1. Preliminaries.
@@ A.3. Construction of $\phi_{t}$
Definition A.23 (Well-Order).
@@
Since $A$ is countable, we can inject $A$ into $\mathbb{N}$. By listing elements of $A$ in the order given by their images under this injection, we get a well‐order $<_{A}$ on $A$. Note that for any non-empty subset $B\subseteq A$, the image of $B$ in $\mathbb{N}$ has a least element, whose preimage is the least element of $B$ under $<_{A}$. We will later use the following remark when constructing $\xi_{t}$ and $\phi_{t}$.
Remark A.24.
By the definition of $<_{A}$, each element of $A$ is either the least element of $A$ or has an immediate predecessor in $A$.
The query set $Q$ is not necessarily countable. However, by the well‐ordering theorem (which is equivalent to the axiom of choice), there exists a well‐order $<_{Q}$ on $Q$, regardless of its cardinality.
Definition A.25 (Lexicographic Order $<_{t}$ on $(Q\times A)^{t}$).
For $t\in\mathbb{N}\cup\{0\}$, we define an order $<_{t}$ on $(Q\times A)^{t}$ by lexicographic comparison: For any two distinct tuples $(q_{1},\dots,a_{t})$ and $(q_{1}^{\prime},\dots,a_{t}^{\prime})$ in $(Q\times A)^{t}$, let $i$ be the smallest index such that $(q_{i},a_{i})\neq(q_{i}^{\prime},a_{i}^{\prime})$. Then:
$(q_{1},\dots,a_{t})<_{t}(q_{1}^{\prime},\dots,a_{t}^{\prime})\text{ if and only if }\left[q_{i}<_{Q}q_{i}^{\prime}\right]\text{ or }\left[q_{i}=q_{i}^{\prime}\text{ and }a_{i}<_{A}a_{i}^{\prime}\right]$
Since $<_{A}$ and $<_{Q}$ are well-orders, $<_{t}$ is also a well-order: $<_{t}$ is clearly a total order. We will show that every non-empty subset $B\subseteq(Q\times A)^{t}$ has a least element. Define
$B_{1}^{Q}$ =$=\{q\in Q\mid\exists a_{1},\dots,a_{t}\in A,\exists q_{2},\dots,q_{t}\in Q\text{ such that }(q,a_{1},q_{2},\dots,a_{t})\in B\}\text{ and }$
$B_{1}^{A}$ =$=\{a\in A\mid\exists a_{2},\dots,a_{t}\in A,\exists q_{2},\dots,q_{t}\in Q\text{ such that }(q_{1}^{*},a,q_{2},\dots,a_{t})\in B\}.$
Since $B$ is non-empty, $B_{1}^{Q}$ is also non-empty. By the well-ordering of $Q$, $B_{1}^{Q}$ has a least element $q_{1}^{*}$. Similarly, since $B_{1}^{A}$ is non-empty, it has a least element $a_{1}^{*}$.
Again, since $B_{2}^{Q}$ is non-empty, it has a least element $q_{2}^{*}$. Repeating this process $t$ times, we obtain the least element $(q_{1}^{*},\dots,a_{t}^{*})$ of $B$.
Lemma A.26 (Transfinite Recursion).
@@
Let $f$ be a function recursively defined over $(Q\times A)^{t}$ according to the well-order $<_{t}$. If for each sequence $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the value of $f(q_{1},\dots,a_{t})$ is uniquely determined by the values $f(q_{1}^{\prime},\dots,a_{t}^{\prime})$ for every $(q_{1}^{\prime},\dots,a_{t}^{\prime})<_{t}(q_{1},\dots,a_{t})$, then $f$ exists and is unique.
In later sections, we define functions $f$ recursively over $(Q\times A)^{t}$ using the well-order $<_{t}$. Specifically, we determine $f(q_{1},\dots,a_{t})$ based on the values $f(q_{1},\dots,a_{t-1},q_{t},a)$ for every $a<_{A}a_{t}$. By the definition of $<_{t}$, we have:
$(q_{1},\dots,a_{t-1},q_{t},a)<_{t}(q_{1},\dots,a_{t}).$
Finally, we define two key sets that will be used in our constructions: for any $a_{t}\in A$,
$\Gamma_{<a_{t}}=\{a\in A\mid a<_{A}a_{t}\},$
and
$\Gamma_{>a_{t}}=\{a\in A\mid a_{t}<_{A}a\}.$
A.3.2. Defining $\phi_{t}$ and showing its existence.
@@
We define the function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ recursively for each $(q_{1},\dots,a_{t})$ based on the $<_{t}$ order. For each $(q_{1},\dots,a_{t})$, we determine $\phi_{t}(q_{1},\dots,a_{t})$ by solving an optimization problem with the variables $\phi_{t}^{0}(q_{1},\dots,a_{t})$ and $\phi_{t}^{1}(q_{1},\dots,a_{t})$. The constraints for this problem involve the known functions $\mu_{t}$, $\phi_{t-1}$, $\mathrm{L}_{t}$, and the known values $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a)$ for every $a\in\Gamma_{<a_{t}}$.
We define $\phi_{t}^{0}(q_{1},\dots,a_{t})$ and $\phi_{t}^{1}(q_{1},\dots,a_{t})$ as the largest possible real numbers in $[0,1]$ satisfying the following constraints:
  1. ($\phi$1)
@@
    $\sum_{a\in\Gamma_{<a_{t}}}$ $\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)+\phi_{t}^{b}(q_{1},\dots,a_{t})+\sum_{a\in\Gamma_{>a_{t}}}{\frac{1}{\delta}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
  3. ($\phi$3)
@@
    $\delta\phi_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\delta\phi_{t}^{1-b}(q_{1},\dots,a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
Although $\Gamma_{<a_{t}}$ is always finite, $\Gamma_{>a_{t}}$ could be infinite. Thus, to ensure ($\phi$2) is well-defined, we need to show:
Claim A.27.
For any $t\in[T]$, $b\in\{0,1\}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, the summation $\sum_{a\in\Gamma_{>a_{t}}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ exists.
@@
By Lemma A.18,
$\sum\limits_{a\in\Gamma_{>a_{t}}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\sum\limits_{a\in A}\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)=\mu_{t-1}^{b}(q_{1},\dots,a_{t-1})\leq 1.$
Moreover each summand $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ is non-negative, implying that the summation is within the range $[0,1]$, and thus it exists. ∎
Claim A.28.

### v3

@@ 8. Reduction to Interactive Randomized Response
implying that $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is a PMF.
8.2.4. Proof of Lemma 5.6.
By definition of $\mathrm{IRR}_{\epsilon,\delta}$ and $\mathcal{P}$, for every $t\in\mathbb{N}$ and every transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the probability of $\mathcal{P}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ returning the answers $a_{1},\dots,a_{t}$ to the queries $q_{1},\dots,q_{t}$ is
@@
Thus, the views of any adversary interacting with $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ are identically distributed, implying that these mechanisms are equivalent.
Furthermore, for every query sequence $(q_{1},\dots,q_{k})$, if the answer distributions of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to $(q_{1},\dots,q_{k})$ are identical, then by Lemma 8.10, for every $t\leq k$ and every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$, we have
$\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t}).$
Therefore, by condition (IV), for every $t\leq k$ and every $(a_{1},\dots,a_{t})\in A^{t}$,
$\phi_{t}^{0}(q_{1},\dots,a_{t})=\phi_{t}^{1}(q_{1},\dots,a_{t}).$
Consequently,
$\psi_{t}^{0}(q_{1},\dots,a_{t})=\psi_{t}^{1}(q_{1},\dots,a_{t}).$
By the definitions of $\mathcal{F}_{(\top,0)}$, $\mathcal{F}_{(\top,1)}$, $\mathcal{F}_{(\bot,0)}$, and $\mathcal{F}_{(\bot,1)}$, this implies that for every $t\leq k$ and answer sequence $(a_{1},\dots,a_{t-1})$, the PMF associated with $(q_{1},\dots,a_{t-1},q_{t})$ is identical in both $\mathcal{F}_{(\bot,0)}$ and $\mathcal{F}_{(\bot,1)}$, and likewise in $\mathcal{F}_{(\top,0)}$ and $\mathcal{F}_{(\top,1)}$. Hence, by the construction of $\mathcal{P}$, this mechanism interacts with $\mathrm{IRR}_{\epsilon,\delta}$ only once. Moreover, by design, $\mathcal{P}$ never interacts with $\mathrm{IRR}_{\epsilon,\delta}$ more than twice.
8.3. Construction of $\phi_{t}$
For each $q_{1},\dots,a_{t-1},q_{t}$, Lyu [19] initializes the values of $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ with the ones from $\mathrm{L}_{t,T}(q_{1},\dots,a_{t-1},q_{t},\cdot)$. (Removing the upper bound assumption, they are initialized by the values of $\mathrm{L}_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$.) Then, they iterate over all possible answers $a_{t}\in A$ in a fixed order and gradually increase the value of $\phi_{t}(q_{1},\dots,a_{t})$ to satisfy conditions (II) and (III), while ensuring that condition (I) remains valid. This iterative procedure requires $A$ to be finite; otherwise, the construction algorithm does not terminate. Moreover, the fixed iteration order can lead to an imbalance in which earlier values grow significantly larger than those that come later in the order. In our new construction, we control and balance this increase.
Intuitively, we introduce a temporary upper bound and increase the values of the function $\mathrm{L}_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ to reach the minimum of the original upper bounds and the new temporary ones. This results in an intermediate function, denoted by $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$. Then, we remove the temporary upper bound and continue increasing the values of $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ (if needed) to meet the original upper bounds. The final function is the desired $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$. We will show that if $\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t})$ for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$, then the functions $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ and $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ are identical. We set the temporary upper bound in a way that this implies the condition (IV) holds.
Formally, we inductively construct $\phi_{t}$ for each $t\in\mathbb{N}$, assuming that $\phi_{t-1}$ exists and satisfies conditions (I), (II) and (IV). We then show that $\phi_{t}$ satisfies all conditions from (I) to (IV). For the base case $t=0$, we define $\phi_{0}()=(1,1)$. By Corollary 8.15, this choice satisfies condition (I). Moreover, for $\phi_{0}^{0}()=\phi_{0}^{1}()=1$ and $\mu_{0}^{0}()=\mu_{0}^{1}()=1$, condition (II) is $\delta\cdot 1-e^{-\epsilon}\cdot\delta\cdot 1\leq 1-e^{-\epsilon}\cdot 1$, which is true because $\delta\leq 1$. Furthermore, since $\phi_{0}^{0}()=\phi_{0}^{1}()=1\geq\min\{1,\delta+\delta\}$, condition (IV) is also satisfied.
Remark 8.21.
In the rest of this section, we assume $\phi_{t-1}$ exists and satisfies conditions (I), (II) and (IV).
Section 8.3.1 outlines the preliminaries required for defining $\xi_{t}$ and $\phi_{t}$. In Section 8.3.2, we define the intermediate function $\xi_{t}$. Section 8.3.3 presents the new construction of $\phi_{t}$, and finally, in Section 8.3.4, we prove that $\phi_{t}$ satisfies conditions (I) to (IV).
8.3.1. Preliminaries.
@@ 8.3. Construction of $\phi_{t}$
Definition 8.22 (Well-Order).
@@
Since $A$ is countable, $A$ can be injected into $\mathbb{N}$. Listing elements of $A$ in the order given by their images under this injection gives a well‐order $_{A}$ on $A$. Note that for any non-empty subset $B\subseteq A$, the image of $B$ in $\mathbb{N}$ has a least element, whose preimage is the least element of $B$ under $_{A}$. We will later use the following remark when constructing $\phi_{t}$.
Remark 8.23.
By the definition of $_{A}$, each element of $A$ is either the least element of $A$ or has an immediate predecessor in $A$.
Definition 8.24 ($\Gamma_{<a}$ and $\Gamma_{>a}$).
For any answer $a\in A$, $\Gamma_{<a}=\{a^{\prime}\in A\mid a^{\prime}<_{A}a\}$ and $\Gamma_{>a}=\{a^{\prime}\in A\mid a<_{A}a^{\prime}\}$.
Lemma 8.25 (Transfinite Recursion [22]).
@@
Let $f$ be a function recursively defined over $A$ according to the well-order $_{A}$. If for each answer $a\in A$, the value of $f(a)$ is uniquely determined by the values $f(a^{\prime})$ for every $a^{\prime}\in\Gamma_{<a}$, then $f$ exists and is unique.
In the next section, we construct functions $f:A\to[0,1]^{2}$ and apply Lemma 8.25 to prove that they are well-defined. To apply this lemma, we must ensure that each value $f(a)$ is uniquely determined based on the previously defined values $f(a^{\prime})$ for all $a^{\prime}\in\Gamma_{<a}$. Thus, to define $f(a)$, we first select a subset $\mathcal{C}\subseteq[0,1]^{2}$ of desired values satisfying certain properties derived from the values of $f(a^{\prime})$ for $a^{\prime}\in\Gamma_{<a}$. We then set $f(a)$ to the lexicographically maximal element of $\mathcal{C}$. The following definitions and lemma formalizes this and guarantee that $f(a)$ is uniquely defined in this way.
Definition 8.26 (Maximal Pair).
Let $\mathcal{C}\subseteq[0,1]^{2}$. A pair $(x_{0},x_{1})\in\mathcal{C}$ is said to be maximal in $\mathcal{C}$ if there does not exist a pair $(y_{0},y_{1})\in\mathcal{C}$ such that $(x_{0},x_{1})\neq(y_{0},y_{1})$, $x_{0}\leq y_{0}$, and $x_{1}\leq y_{1}$.
Definition 8.27 (Lexicographic Order $_{lex}$).
The relation $_{lex}$ is a binary ordering on $0,1]^{2}$ defined as follows: for any $(r_{1},r_{2}),(r_{3},r_{4})\in[0,1]^{2}$, we have $(r_{1},r_{2})<_{lex}(r_{3},r_{4})$ if and only if $r_{1}\leq r_{3}$, or $r_{1}=r_{3}$ and $r_{2}\leq r_{4}$.
Lemma 8.28.
Let $\mathcal{C}$ be a non-empty compact subset of $0,1]^{2}$.
  1. (a)
    $(x_{0},x_{1})<_{lex}(x_{0}^{*},x_{1}^{*})$ for all $(x_{0},x_{1})\in\mathcal{C}$.
  2. (b)
    $(x_{0}^{*},x_{1}^{*})$ is maximal in $\mathcal{C}$.
Proof.
(a): Define $x_{0}^{*}=\sup\{x_{0}\mid(x_{0},x_{1})\in\mathcal{C}\}$. Since $\mathcal{C}$ is non-empty, the set $\{x_{0}\mid(x_{0},x_{1})\in\mathcal{C}\}$ is non-empty. Also, by definition, $x_{0}^{*}\leq 1$. As we are taking a continuous projection $(x_{0},x_{1})\to x_{0}$ of a compact set $\mathcal{C}$, this supremum is actually a maximum, so there exists $(x_{0},x_{1})\in\mathcal{C}$ such that $x_{0}=x_{0}^{*}$.
Define $x_{1}^{*}=\sup\{x_{1}\mid(x_{0}^{*},x_{1})\in\mathcal{C}\}$. Again, the set $\{x_{1}\mid(x_{0}^{*},x_{1})\in\mathcal{C}\}$ is non-empty since there exists $(x_{0},x_{1})\in\mathcal{C}$ with $x_{0}=x_{0}^{*}$. By the same compactness argument, this supremum is also attained. Hence, $(x_{0}^{*},x_{1}^{*})\in\mathcal{C}$. By definition, $(x_{0},x_{1})<_{lex}(x_{0}^{*},x_{1}^{*})$ for every $(x_{0},x_{1})\in\mathcal{C}_{q_{1},\dots,a_{t}}$, which trivially implies that $(x_{0}^{*},x_{1}^{*})$ is the only pair in $\mathcal{C}$ satisfying this property.
(b): Let $(x_{0},x_{1})$ be a pair in $\mathcal{C}$ satisfying $x_{0}\geq x_{0}^{*}$ and $x_{1}\geq x_{1}^{*}$. Since $x_{0}\geq x_{0}^{*}$ and $(x_{0},x_{1})\in\mathcal{C}$, by definition of $x_{0}^{*}$, we have $x_{0}^{*}=x_{0}$. Since $x_{1}\geq x_{1}^{*}$, $x_{0}^{*}=x_{0}$, and $(x_{0},x_{1})\in\mathcal{C}$, by definition of $x_{1}^{*}$, we have $x_{1}^{*}=x_{1}$. Therefore, $(x_{0},x_{1})=(x_{0}^{*},x_{1}^{*})$, implying that $(x_{0}^{*},x_{1}^{*})$ is maximal. ∎
8.3.2. Defining $\xi_{t}$ and showing its existence.
@@
To define the function $\xi_{t}:(Q\times A)^{t}\to[0,1]^{2}$, we fix $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$ and recursively define $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ for each answer $a_{t}\in A$ according to the order $_{A}$. We define $\xi_{t}(q_{1},\dots,a_{t})=(x_{0}^{*},x_{1}^{*})$, where $x_{0}^{*}$ and $x_{1}^{*}$ are uniquely chosen from the maximal feasible solutions of a constraint satisfaction problem with two variables $x_{0}$ and $x_{1}$. The constraints for this problem involve the known functions $\mu_{t}$, $\phi_{t-1}$, $\mathrm{L}_{t}$, and the known values $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},a)$ for every $a\in\Gamma_{<a_{t}}$.
Consider the following constraints for the variables $x_{0}$ and $x_{1}$:
  1. ($\xi$.1)
@@
    $\sum_{a\in\Gamma_{<a_{t}}}$ $\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)+x_{b}+\sum_{a\in\Gamma_{>a_{t}}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
  3. ($\xi$.3)
@@
    $x_{b}-e^{-\epsilon}x_{1-b}\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
  4. ($\xi$.4)
    For each $b\in\{0,1\}$,
    $x_{b}\leq\min\{\mu_{t}^{b}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$
Let the set $\mathcal{C}_{q_{1},\dots,a_{t}}$ denote the set of feasible solutions $(x_{0},x_{1})$ that satisfy all the constraints above. Each constraint is a non-strict inequality of the form $h_{i}(x_{0},x_{1})\geq 0$, where each $h_{i}$ is a continuous function. Since the solution set of a single non-strict inequality defined by a continuous function is closed, and the intersection of finitely many closed sets is also closed, the set $\mathcal{C}_{q_{1},\dots,a_{t}}$ is closed. Furthermore, as it is a subset of the compact set $0,1]^{2}$, it is itself compact. If $\mathcal{C}_{q_{1},\dots,a_{t}}$ is nonempty, then by Lemma 8.28, the lexicographically maximum pair $(x_{0}^{*},x_{1}^{*})$ of $\mathcal{C}_{q_{1},\dots,a_{t}}$ exists, belongs to $\mathcal{C}_{q_{1},\dots,a_{t}}$, and is maximal in $\mathcal{C}_{q_{1},\dots,a_{t}}$. We define $\xi_{t}(q_{1},\dots,a_{t})=(x_{0}^{*},x_{1}^{*})$.
Claim 8.29.
The function $\xi_{t}$ is well-defined and exists.
@@
The proof consists of two steps: (i) proving that $\mathcal{C}_{q_{1},\dots,a_{t}}$ is non-empty, and (ii) applying the transfinite recursion theorem (Lemma 8.25) and concluding that $\xi_{t}$ exists.
(i): Fix any $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$. By Remark 8.23, each answer in $A$ is either the smallest element of $A$ or has an immediate predecessor. We will inductively show that for every $a_{t}\in A$, the values $\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})$ and $\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})$ satisfy the constraints for defining $\xi_{t}(q_{1},\dots,a_{t})$, and consequently, $\mathcal{C}_{q_{1},\dots,a_{t}}$ is not empty.
Trivially, $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})$ satisfies Constraint ($\xi$.1). Moreover, by Corollary 8.16, it satisfies Constraint ($\xi$.3). To prove Constraint ($\xi$.2) is satisfied, we need to show that
(2) $\sum_{a\in\Gamma_{<a_{t}}}\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)$ $+\sum_{a\in\Gamma_{>a_{t}}\cup\{a_{t}\}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\delta\phi_{t-1}^{b}(q_{1},\dots,a_{t-1}),$
We prove this by induction on $a_{t}\in A$:
We know that $\phi_{t-1}(q_{1},\dots,a_{t-1})$ satisfies condition (I). (see Remark 8.21.) Thus, by Lemma 8.20,
finishing the proof for the base case.
Inductive Step: Let $a^{*}$ denote the immediate predecessor of $a_{t}$. By inductive hypothesis, $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ exists and satisfies Constraint ($\xi$.3). Since $\Gamma_{<a_{t}}=\Gamma_{<a^{*}}\cup\{a^{*}\}$, Constraints ($\xi$.2) for defining $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},a^{*})$ is identical to (2). Thus, by inductive hypothesis, inequality (2) holds, completing the induction.
It remains to prove that Constraint ($\xi$.4) is satisfied. By Lemma 8.17, $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})\leq\mu_{t}^{b}(q_{1},\dots,a_{t})$. Also, as $\mathrm{L}_{t}^{1-b}$ is non-negative, $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})\leq\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})$. Thus,
$\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t})\leq\min\{\mu_{t}^{b}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\},$
and Constraint ($\xi$.4) is satisfied.
(ii): For every $(q_{1},\dots,a_{t-1},q_{t})$, the values of the function $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ is uniquely determined by the preceding values of this function in the well-order $_{A}$. Thus, by the transfinite recursion theorem, the function $\xi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ exists and is uniquely defined. Consequently, the function $\xi_{t}$ exists and is uniquely defined. ∎
8.3.3. Defining $\phi_{t}$ and showing its existence.
To define the function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$, we fix $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$ and recursively define $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ for each answer $a_{t}\in A$ based on the $_{A}$ order. Similar to the definition of $\xi_{t}$, for each $(q_{1},\dots,a_{t})$, we determine $\phi_{t}(q_{1},\dots,a_{t})$ using a set of constraints. These constraints involve the known functions $\xi_{t}$, $\mu_{t}$, $\phi_{t-1}$, $\mathrm{L}_{t}$, and the known values $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},a)$ for every $a\in\Gamma_{<a_{t}}$.
Consider the following set of constraints for the variables $x_{0}$ and $x_{1}$:
  1. ($\phi$.1)
    For each $b\in\{0,1\}$,
    $\delta x_{b}\geq\xi_{t}^{b}(q_{1},\dots,a_{t}).$
  2. ($\phi$.2)
    For each $b\in\{0,1\}$,
    $\sum_{a\in\Gamma_{<a_{t}}}$ $\phi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)+x_{b}+\sum_{a\in\Gamma_{>a_{t}}}{\frac{1}{\delta}}\xi_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
  3. ($\phi$.3)
    For each $b\in\{0,1\}$,
    $\delta x_{b}-e^{-\epsilon}\delta x_{1-b}\leq\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$
Let $\mathcal{C}_{q_{1},\dots,a_{t}}$ denote the set of feasible solutions $(x_{0},x_{1})$ satisfying the constraints above. As in Section 8.3.2, one can show that $\mathcal{C}_{q_{1},\dots,a_{t}}$ is compact. Therefore, if $\mathcal{C}_{q_{1},\dots,a_{t}}$ is nonempty, then by Lemma 8.28, the lexicographically maximum pair $(x_{0}^{*},x_{1}^{*})$ of $\mathcal{C}_{q_{1},\dots,a_{t}}$ exists, belongs to $\mathcal{C}_{q_{1},\dots,a_{t}}$, and is maximal in $\mathcal{C}_{q_{1},\dots,a_{t}}$.
Define $\mathcal{G}_{q_{1},\dots,a_{t}}\subseteq\mathcal{C}_{q_{1},\dots,a_{t}}$ to be the set of all maximal elements in $\mathcal{C}_{q_{1},\dots,a_{t}}$. By the definition of maximality, this set either contains no pair of the form $(z,z)$ with identical components, or it contains exactly one such pair, denoted $(z^{*},z^{*})$. We define $\phi_{t}(q_{1},\dots,a_{t})=(z^{*},z^{*})$ if there exists $(z^{*},z^{*})\in\mathcal{G}_{q_{1},\dots,a_{t}}$ and $\phi_{t}(q_{1},\dots,a_{t})=(x_{0}^{*},x_{1}^{*})$ otherwise. In either case, $\phi_{t}(q_{1},\dots,a_{t})$ is a maximal solution for conditions ($\phi$.1) to ($\phi$.3).
Claim 8.30.